For the following situation, add or subtract the vectors to find the magnitude of the resultant.

Emma rides her bike 5km north, stops for lunch, and then rides 4km north.

Magnitude =   
Answer #1

We consider 5 km North 4 km Worth Resultant, North as ty axis. Ã = 5 km } ★ 7 = 4 km } Ř = Ã +3 = 59 +49 - 9 km î Magnitude =